what is the sum of the first 40 terms whose general term is An=5n+10


1
Expert's answer
2021-12-27T18:33:18-0500

"A1 = 5\u00d71 + 10 = 15,"

"A40 = 5\u00d740 + 10 = 210."

The sum of the first 40 terms is:

"S40 = 40\u00d7(15 + 210)\/2 = 4,500."